OpenStudy (anonymous):

PLEASE HELP ME!!!!! I NEED HELP REALLY BAD!!!! You are making poster board boundary markers in the shape of square pyramids for a relay race. If you are including a 10 -inch square as the bottom and you want the slant height to be 12 inches, how much poster board do you need?

OpenStudy (anonymous):

340 inches

OpenStudy (anonymous):

oh ok thank you!!!!!!! :)

OpenStudy (btaylor):

 you have 1 square, area 100. Then each triangle has area (0.5 x 10 x 12)=60, and there are four. so 100 + 4(60)= 100+240 = 340